OpenStudy (anonymous):

One number is 3 times another. If the sum of their reciprocals is 92 , find the two numbers.

OpenStudy (anonymous):

if one is \(x\) the the other is \(3x\) the sum of their reciprocals is \[\frac{1}{x}+\frac{1}{3x}=\frac{4}{3x}\]

OpenStudy (anonymous):

set \[\frac{4}{3x}=92\] and solve for \(x\)